OpenStudy (anonymous):

Part 1: Show the expanded form and simplification of x^4/x^2 . Explain in your own words how you can simplify x^4/x^2 without having to write the expanded form. Part 2: Create your own fraction with like bases, coefficients, and show its simplification.

OpenStudy (anonymous):

i know i fixed it

OpenStudy (anonymous):

\[\frac{ x^4 }{ x^2 }\] When dividing common terms, SUBTRACT the exponents. So your answer should be x^2.

OpenStudy (anonymous):

thanks but can you show me how exactly you did that

OpenStudy (anonymous):

It's just a rule of algebra... If you want you can write it like this \[\frac{ x^4 }{ x^2 }=\frac{ x^2*x^2 }{ x^2 }=x^2\] But that's kinda stupid. When dividing powers, subtract them.

OpenStudy (anonymous):

ok thank you